# A new criterion for interpreting acoustic emission damage signals in condition monitoring based on the distribution of cluster onsets

### v1

@@ Abstract
Structural Health Monitoring (SHM) relies on non-destructive techniques such as Acoustic Emission (AE) which provide a large amount of data over the life of the systems. The analysis of these data is often based on clustering in order to get insights about damage evolution. In order to evaluate clustering results, current approaches include Clustering Validity Indices (CVI) which favor compact and separable clusters. However, these shape-based criteria are not specific to AE data and SHM. This paper proposes a new approach based on the sequentiality of clusters onsets. For monitoring purposes, onsets indicate when potential damage occurs for the first time and allows to detect the inititation of the defects. The proposed CVI relies on the Kullback-Leibler divergence and enables to incorporate prior on damage onsets when available. Three experiments on real-world data sets demonstrate the relevance of the proposed approach. The first benchmark concerns the detection of the loosening of bolted plates under vibration. The proposed onset-based CVI outperforms the standard approach in terms of both cluster quality and accuracy in detecting changes in loosening. The second application involves micro-drilling of hard materials using Electrical Discharge Machining. In this industrial application, it is demonstrated that the proposed CVI can be used to evaluate the electrode progression until the reference depth which is essential to ensure structural integrity. Lastly, the third application is about the damage monitoring in a composite/metal hybrid joint structure. As an important result, the timeline of clusters generated by the proposed CVI is used to draw a scenario that accounts for the occurrence of slippage leading to a critical failure.
[…]
${}^{1*}$, Université de Franche-Comté, CNRS, institut FEMTO-ST, F-25000, Besançon, France
${}^{2}$ristol University, Department of Aerospace Engineering, United Kingdom
${}^{3}$niversité de Franche-Comté, CNRS, institut FEMTO-ST, F-25000, Besançon, France
${}^{4}$niversité de Franche-Comté, Institut FEMTO-ST, F-25000, Besançon, France
${}^{5}$epartment of Materials, The University of Manchester, Manchester, UK & Aerospace Research Institute, The University of Manchester, Manchester, UK
${}^{6}$i-Composites Lab, Department of Materials Science and Engineering & Department of Mechanical and Aerospace Engineering, Monash University, Clayton, VIC, Australia
$*{}^{*}$emmanuel.ramasso@femto-st.fr
@@ 1 Introduction
Acoustic emission (AE) is a non-destructive technique developed to monitor materials during mechanical tests. The data are collected through highly sensitive sensors able to detect subnanometric displacements at the surface of a material. These displacements occur due to the propagation of stress waves generated by the formation of damage within the material/structure. Therefore, the AE technique is widely used as inputs of expert systems to get insights about the behaviour of a material or a structure under stress [1]. Due to the lack of knowledge concerning the precise source of each AE signal, expert systems take the form of unsupervised learning workflows [2, 3] based on three main steps in order to interpret AE data (Figure 1).
@@
During experimental campaigns, AE data are collected in the form of a data stream acquired between 1 and 10 MHz. The stream is made up of AE transients mixed with noise. The first processing step includes wave-picking in order to detect discrete AE signals (also called AE hits) that are deemed to be relevant (i.e. not noise) for interpreting the material behavior [4, 5, 6, 7]. Each AE signal is traditionally transformed into a $d$-dimensional feature vector that represents a basis with the same number of dimensions and axes for all signals. Some dozens of features are generally extracted [8] due to a lack of knowledge about the optimal subset of features to be used. The features are represented by a $N\times d$ matrix that feeds into the clustering algorithm applied in the second step. The goal of clustering is to create $K$ groups of feature vectors called clusters. The clustering step involves hyperparameters (the algorithm’s parametrisation), which play a role in the convergence but can not be optimised through the objective function used to find the clusters. Two critical hyperparameters are generally involved in AE data clustering: the choice of the features and the choice of the number of clusters.
In the standard approach for AE data clustering, these hyperparameters are generally set according to an algorithm detailed in A which includes the third step, called validation. This step makes use of a criterion (or a combination of criteria) to evaluate the quality of the clustering result according to the hyperparameters. In pattern recognition, these criteria are also called Clustering Validity Indices (CVI) [9]. Following [10, 2], several indices (line 14) can be computed for all parametrisations ($k\cdot C_{n}^{p}$ for subsets of $p$ features from $n$ and considering $k$ different numbers of clusters). Then a majority voting scheme, relying on some weighting provided by the end-user, selects the best number of clusters and the best subset of features (lines 19-24).
“Quality” in clustering is a subjective notion. For decades, researchers interested in AE data clustering have focused on finding "natural" clusters. A natural cluster was defined in 1968 by J. W. Carmichael, R. S. Julius [11] as "(1) a relatively densely populated regions of the space, and (2) surrounded by continuous and relatively empty regions of the space". Given this definition related to the shape and compactness of clusters in the feature space, the clustering of AE data sets collected during mechanical tests is generally posed as finding a set of hyperparameters such that natural clusters are obtained. For that, internal CVI are almost exclusively used to evaluate to what extent a set of clusters is natural for a given AE data set. An internal CVI takes the features and the parameters of the clusters (e.g. centers) as inputs, in addition to some hyperparameters involved in the clustering method (for example, the type of distance between feature vectors). It then generates a real value that reflects the quality of the clustering result. In AE literature, Davies-Bouldin [12] and Silhouette criteria [13] are often used. By focusing on clusters’ shape, these CVI foster clustering results made of compact and well-separated clusters.
As an alternative to focusing on clusters’ shape for the validation of clustering results, several publications were recently focused on the development of algorithms for finding relevant clusters in terms of sequence, timeline or chronology. For example, a visualization of AE events in geoscience was suggested in [14] where the authors determined the 3D location through time of each AE event, similar to locating the focus of an earthquake. Similarly, an approach integrating spectral analysis of AE waveforms was proposed in [15, 16] aiming at identifying friction and wear mechanisms during degradation of lubricants and contacting surfaces in tribological tests. The authors focused on the chronology of damages and their approach based on supervised learning was shown to simplify the presentation of AE results for quick interpretation. Some other authors did not use clustering but cumulated features in order to visually interpret AE signals and to evaluate damage progression. Cumulated features generally represents a monoparametric interpretation of AE signals and is a different approach compared to clustering-based analysis which is multivariate [17]. Another drawback of cumulated features is their physical meaning. Moreover, the cumulated values can hide some AE signals which are of low amplitude but frequent and can represent the initiation of late damages.
The idea behind these methods is that damage mechanisms do not usually start simultaneously but according to a timeline: One is often the precursor of the others and damage follows a progression, as known from physics [18], [19, Chap. 6]. In terms of data science dedicated to AE data, to our knowledge, there is no formal definition of a CVI that allows to select the best parametrisation of a clustering method using onsets of clusters. Assuming that materials under stress degrade progressively, previous publications shown that such a CVI can be useful for onset-based interpretation of AE. A focus on the timeline of clusters can provide insights into how damages progressively occur and develop. It can inform about the dynamic process involved in the generation of AE signals, which may be relevant to better understand the material behavior and can be used for structural health monitoring purpose [20, 21, 22].
@@
Therefore, our contribution is the formulation of a new CVI for interpreting AE signals based on the distribution of clusters onsets and its validation on three applications. In order to illustrate the positioning of our contribution, Figure 2 depicts the fundamental difference between shape-based and onset-based approaches for the evaluation of AE data clustering results. With the former approach (top subfigures), clusters may be compact and separated in a low-dimensional feature space, but are not ensured to lead to relevant onsets. For example several of them can start very early and almost simultaneously in the data. For monitoring purposes, the latter approach aims at finding subsets of features that point out onsets distributed through the horizontal axis (load or time) which is more helpful to interpret the dynamic behavior of the material or structure under test (bottom subfigures).
(a) "Selection" (left) is based on clusters shape in the feature space in low-dimensional space for visualization purpose. "Representation" (right) of clusters of the timeline provides onsets which start too early.
(b) "Selection" (right) is based on clusters onsets. The timeline is relevant for monitoring. The "representation" (left) in low-dimension (2D in general) of the clusters can be useless because the clusters are compact and separated only in higher dimensions (used in clustering). The compactness and separation of clusters is ensured during clustering.
@@
Section 2 introduces the proposed criterion for selecting the best parametrisation. It is evaluated in Section 3 using a benchmark obtained from a real structure. This benchmark allows us to quantify whether the criterion drives the end-user to a better interpretation compared to standard approaches.
2 Sorting subsets of features with respect to onsets
The AE data are supposed to be made of $N$ $d$-dimensional feature vectors $\boldsymbol{x}_{n}\in\Re^{d}$. Given a subset of features $S\subset\Re^{d}$ and a number of clusters $K$, a partition $P$ is obtained by a clustering method. $P$ is represented by a vector of length $N$, where each element $P(n)\in\{1,2\dots K\}$ represents the cluster number assigned to each AE signal. For fuzzy clustering methods, $P(n)$ is the cluster with the largest membership degree.
A cluster’s onset is defined as the first occurrence of the cluster. For example, for cluster $k$, the onset is the first index in $P$ where the $k$-th cluster appears. Onsets can be represented on an axis related to time, load or any physical quantities related to the degradation of the material. Illustrations can be found on the right hand-side of Figure 2 and on Figure 3.
Figure 3: Illustration used in examples in Section 2: Onsets (triangles) are defined as the first occurrence of each cluster. The horizontal axis can be any monotically increasing information regarding the evolution of the test like, for example, time or a damage variable. In-between onsets, several clusters, that previously occurred, can keep growing. Reference to composite damages is only for illustration. Top figure: No signal appears from the beginning to the first onset so this part is removed and the starting time shifted accordingly, which leads to the bottom figure.
Let’s define $t_{k}\in[0,T]$ as the first occurrence of cluster $k$ in a data set (with the first onset starting at $t_{1}=0$). Using onsets, we compute the normalised difference between two consecutive onsets as
@@ 2 Sorting subsets of features with respect to onsets
Example 1.
[…]
2.1 Case 1: A priori on onsets is available
Let’s first assume that the end-user can provide the expected onsets (composed of $K$ values). This prior onsets can, for example, be provided by a physics-based model able to predict the onset of each damage or cluster. In that case, we have a piece of information about onsets that should be taken into account to select the clustering parametrisation. These prexisting onsets can be represented as a distribution, say ${\mathbf{p}}_{\textrm{true}}$, following the same reasoning as before. The problem is now to compare ${\mathbf{p}}_{\textrm{true}}$ (the prior distribution) and ${\mathbf{p}}_{\textrm{est}}$ (the estimated distribution).
[…]
@@ 2.1 Case 1: A priori on onsets is available
In this expression, ${\mathbf{p}}_{\textrm{est}}$ is dependent on $S$, the subset of features used for clustering, and $K$ the number of clusters. The KL is thus made of two terms. First, a cross entropy loss
@@
and, second, Shannon’s entropy
@@
2.2 Case 2: Without a priori on onsets
[…]
When using the uniform distribution of onsets for ${\mathbf{p}}_{\textrm{true}}$ in CE (Eq. 4), we get

### v3

# A Novel Criterion for Interpreting Acoustic Emission Damage Signals Based on Cluster Onset Distribution

@@ Abstract
Structural health monitoring (SHM) relies on non-destructive techniques such as acoustic emission (AE) that generate large amounts of data over the lifespan of systems. Clustering methods are used to interpret these data and gain insights into damage progression and mechanisms. Conventional methods for evaluating clustering results utilise clustering validity indices (CVI) that prioritise compact and separable clusters. This paper introduces a novel approach based on the temporal sequence of cluster onsets, indicating the initial appearance of potential damage and allowing for early detection of defect initiation. The proposed CVI is based on the Kullback-Leibler divergence and can incorporate prior information about damage onsets when available.
Three experiments on real-world datasets validate the effectiveness of the proposed method. The first benchmark focuses on detecting the loosening of bolted plates under vibration, where the onset-based CVI outperforms the conventional approach in both cluster quality and the accuracy of bolt loosening detection. The results demonstrate not only superior cluster quality but also unmatched precision in identifying cluster onsets, whether during uniform or accelerated damage growth. The two additional applications stem from industrial contexts. The first focuses on micro-drilling of hard materials using electrical discharge machining, demonstrating, for the first time, that the proposed criterion can effectively retrieve electrode progression to the reference depth, thus validating the setting of the machine to ensure structural integrity. The final application involves damage understanding in a composite/metal hybrid joint structure, where the cluster timeline is used to establish a scenario leading to critical failure due to slippage.
Emmanuel Ramasso1∗, Martin Mbarga Nkogo1, Neha Chandarana2, Gilles Bourbon1, Patrice Le Moal1, Quentin Lefèbvre1, Martial Personeni1, Constantinos Soutis3, Matthieu Gresil4, Sébastien Thibaud1
1∗Université Marie et Louis Pasteur, CNRS, UTBM, institut FEMTO-ST, F-25000, Besançon, France
2Bristol University, Department of Aerospace Engineering, United Kingdom
3Department of Materials, The University of Manchester, Manchester, UK & Aerospace Research Institute, The University of Manchester, Manchester, UK
4i-Composites Lab, Department of Materials Science and Engineering & Department of Mechanical and Aerospace Engineering, Monash University, Clayton, VIC, Australia
emmanuel.ramasso@femto-st.fr
@@ 1 Introduction
Ensuring the availability of equipment and the integrity of structures is a critical concern within the aeronautical and civil engineering industries, both economically and for user safety. This necessitates the monitoring of structures as well as understanding potential damage mechanisms that can indicate the integrity of the structure. Structural health monitoring (SHM) is a research and development field aimed at implementing and deploying solutions to monitor engineering systems such as equipment and structures [18, 20, 33].
During operation, systems and structures are subjected to mechanical, electrical, or chemical stresses, which generate localised micro-deformations. These micro-deformations may be accompanied by the release of energy in the form of elastic waves, known as acoustic emission (AE). When these changes are irreversible and likely to degrade performance, they are referred to as damage. Accumulated micro-damages eventually lead to material failure, rendering the structure unusable, and potentially leading to catastrophic failure [32]. Piezoelectric transducers can be placed onto or embedded in the host structure to record and convert these micro-vibrations into usable electrical signals [20, 9]. The recorded signals contain the signatures of the types of damage from which they have been generated [59, 24, 1, 37].
As a passive technique, it exploits the transient signals emitted during damage occurrence. Therefore, sensors must be continuously monitored or triggered by significant damage. Although the sampling frequency is relatively high (typically 5 to 10 MHz, depending on the sensors), it is possible to detect damage in real time from AE data streams, even during continuous monitoring, thanks to modern computing capabilities. A key challenge in real-time processing is the acoustic emissivity of the monitored material and/or structure; the more emissive the material and/or structure, the greater the number of AE sources it has, leading to more signals that need to be processed, which increases computational costs. To manage this, an amplitude threshold is typically applied to limit the number of transients to be analysed. After distinguishing between noise and meaningful data, various methodologies can be applied to interpret AE data streams. Generally, these methodologies share a common focus on analysing certain statistical properties, although the methods for deriving these properties vary significantly across publications. The relevance of the AE technique has been demonstrated in various applications for decades, in particular for tool wear condition monitoring during in milling, grinding and drilling, as well as for characterisation and monitoring of various materials, such as composite materials and rocks [12, 2, 55].
Correctly interpreting acoustic emission data is of key importance for predictive maintenance in machining processes [19]. However, due to the lack of knowledge concerning the precise source of each AE signal, the interpretation of acoustic emission data remains an open issue.
One approach is sample-based analysis. Samples can be analyzed point by point [42, 40] for wave picking and clustering. Another way is by a sliding window. For example, Martin-Del-Campo et al. [34, 35] developed an anomaly detector where part of the data is labelled as "normal", and the algorithm detects deviations, providing insights into structural integrity of bearings with particle-contaminated lubricant at varying rotational speeds. The method relies on dictionary learning using waveforms matched to a windowed portion of the stream, optimised to produce a sparse and accurate code, thus reducing the measurement data rate by an order of magnitude.
Data analysis can also be performed at the transient level. Transient signals must first be detected using wave selection methods from the literature [29, 41, 31, 43, 24, 28]. Interpretation follows and aims to address challenges such as damage localisation or clustering for damage classification. For example, Chen et al. [11] applied dynamic time warping to measure the similarity of AE signals for source localisation when multiple AE sources are present. Data from thermal-cracking experiments on granite were used to validate the approach. In [46], randomly selected signals are represented by auto-regressive hidden Markov models, subsequently applied to all transients of a dataset to obtain scores indicating goodness-of-fit, which are then used as inputs for clustering algorithms. This method was validated on datasets from carbon fibre composites.
One of the most frequently applied methodologies encountered in the literature relies on three main steps (Figure 1): 1) hit detection - where transient signals are detected based on signal processing tools - and feature extraction where the transient signals are compressed and represented in a common high dimensional space for further analysis, 2) clustering or classification, which either groups features in an unsupervised manner or uses prior labels in a supervised approach to predict signal classes, and 3) evaluation and validation of the groups or classes of signals.
@@
This methodology has been applied, for example, to carbon fibre composite panels for aeronautics [1], high-performance carbon fibre-reinforced thermoset composites in extreme environments [45], bridge and tunnel health assessment [30], railway systems [7], tribology [39, 56], and fatigue crack monitoring [26]. Deep learning approaches, which are supervised methods, follow similar steps but typically automate feature extraction through layers during training [16, 22, 47].
The final step, namely validation, which is central to this work, is more challenging for clustering than for supervised learning due to its subjectivity [8], especially given the lack of precise ground truth in the interpretation of AE data. Various criteria, known as clustering validity indices (CVI), have been proposed in the pattern recognition community [3, 36, 57, 58, 6, 5] to determine whether clusters are meaningful.
For a given application, the challenge for unsupervised interpretation of AE data is the extraction of relevant features that can link the identified clusters to damage mechanisms. This involves answering the following key questions: What are the most relevant features? How many clusters are needed?
An answer to these questions is provided by an algorithm used in the literature detailed in appendix A and proposed by [53]. It uses several CVIs evaluated for different combinations of $k\cdot C_{n}^{p}$ (subsets of $p$ features from $n$, considering $k$ different cluster counts). A majority voting scheme, weighted by the user, is used to select the best number of clusters and subset of features.
The CVI used for AE data interpretation, such as the Davies-Bouldin index [14] or Silhouette index [52], are exclusively focused on clusters’ shape, characterised by their compactness and separability in the feature space. One problem with these CVIs is that the chosen metric (such as Euclidean distance) may differ from those used in the clustering method’s objective function (such as likelihood), potentially leading to misleading interpretations [54, 45]. Sawan et al. [54] showed that acoustic emission data should be divided into more clusters than are expected from the number of active damage mechanisms to prevent misrepresentative grouping. Similarly, the results of several clustering methods were combined in [45] to more accurately map the feature space, resulting in a timeline that reflects a plausible damage scenario. Another issue with classical CVIs is that the timeline provided by compact and separated clusters is generally not related to the expected damage scenario, hence justifying novel approaches to achieve more relevant clustering performance [44, 54, 10, 23, 51, 9, 50]. These works show that damage can be described by a set of clusters, which may or may not correspond to true mechanisms, but the common factor is that one cluster should act as a precursor to others, following a progression.
In previous literature, clustering results are often presented as sequences after applying shape-based CVIs. However, current methods do not assure users that the results will be relevant in terms of the timeline. Therefore, the challenge remains: how can we select features and determine the number of clusters to guarantee a relevant timeline?
Our work aims to address this question by introducing a new validation criterion based on cluster onsets. The main features and advantages of this criterion are as follows:
  • The proposed criterion is the first to formalise the ranking of cluster partitions based on how clusters emerge over time or in relation to a physical parameter. This approach provides a dynamic perspective on clustering behaviour, offering a new way to assess the evolution of clusters beyond static measures. The criterion is innovative and lacks a direct equivalent in the existing literature for selecting the number of clusters and the feature subsets that best explain the data in terms of a timeline.
  • The proposed criterion is valuable for onset-based interpretation of AE signals particularly in cases where materials degrade progressively under stress. Three challenging experiments on real-world datasets demonstrate the relevance of the proposed approach. One of the applications is a benchmark on the detection of the loosening of bolted plates under vibration where the proposed criterion outperforms existing methods. It provides superior results, both quantitatively (in terms of accuracy) and qualitatively (in terms of the positioning of estimated onsets and cumulative occurrences).
@@
To illustrate the positioning of our contribution, Figure 2 highlights the fundamental difference between shape-based and onset-based approaches for evaluating the results of AE data clustering. In the shaped-based approach (Figure 2a), clusters may appear compact and well-separated in a low-dimensional feature space. However, this does not guarantee that they will correspond to relevant onsets. For example several of them can start very early and almost simultaneously when considering the horizontal axis. In contrast, the latter approach focuses on identifying feature subsets that highlight onsets distributed along the horizontal axis (such as load or time). This is more helpful for interpreting the dynamic behaviour of the material or structure under test, as illustrated in Figure 2b.
@@
Section 2 introduces the proposed criterion and presents an algorithm for the implemented method, incorporating the proposed criterion for selecting the best parametrisation and the resulting optimal cluster partition. The relevance of the criterion is evaluated in Section 3 using a benchmark obtained from a real structure.
2 Proposed criterion: Sorting subsets of features with respect to onsets
The AE data consist of $N$ $d$-dimensional feature vectors $\boldsymbol{x}_{n}\in\Re^{d}$. Given a subset of features $S\subset\Re^{d}$ and a number of clusters $K$, a partition $P$ is obtained by applying a clustering method. $P$ is represented by a vector of length $N$, where each element $P(n)\in\{1,2\dots K\}$ represents the cluster number assigned to each AE signal. For fuzzy clustering methods, $P(n)$ is the cluster with the largest membership degree.
A cluster’s onset is defined as the first occurrence of the cluster. For example, for cluster $k$, the onset is the first index in $P$ where the $k$-th cluster appears. Onsets can be represented on an axis related to time, load, or other physical quantities related to the degradation of the material. Illustrations can be found on the right-hand side of Figure 2 and on Figure 3.
Figure 3: Illustration used in examples in Section 2: Onsets (triangles) are defined as the first occurrence of each cluster. The horizontal axis can be any monotonically increasing information regarding the evolution of the test like, for example, time or a damage variable. In-between onsets, several clusters, that previously occurred, can keep growing. Reference to composite damages is only for illustration. Top figure: No signal appears from the beginning to the first onset so this part is removed and the starting time shifted accordingly, which leads to the bottom figure.
Let us define $t_{k}\in[0,T]$ as the first occurrence of cluster $k$ in a dataset. In order to compute the criterion, the onsets are delayed so that the first one starts at $t_{1}=0$ (Figure 3). Then, we compute the normalised difference between two consecutive onsets as
@@ 2 Proposed criterion: Sorting subsets of features with respect to onsets
Example 2.1.
As an example, consider Figure 3. We have three clusters, with onsets positioned at $t_{1}=27,t_{2}=57,t_{3}=82$, and with $T=127$ (horizontal axis is with arbitrary unit; in practice it can be related to time, load, or a physical quantity). We can shift the onsets to $t_{1}=0,t_{2}=30,t_{3}=55$ and $T=100$. Then $\Delta_{1}=30/T$, $\Delta_{2}=25/T$, $\Delta_{3}=45/T$, satisfying Eq. 2. In that case, ${\mathbf{p}}_{\textrm{est}}=[0.30,0.25,0.45]$.
2.1 Case 1: A prior on onsets is available
Let’s first assume that the end-user can provide the expected onsets (composed of $K$ values). These prior onsets can, for example, be provided by a physics-based model able to predict the onset of each damage or cluster. Prior onsets can be represented as a distribution, say ${\mathbf{p}}_{\textrm{true}}$, following the same reasoning as before. The problem is now to compare ${\mathbf{p}}_{\textrm{true}}$ (the prior distribution) and ${\mathbf{p}}_{\textrm{est}}$ (the estimated distribution).
The degree of agreement between estimated onsets and the ground truth as quantified by ${\mathbf{p}}_{\textrm{est}}$ and ${\mathbf{p}}_{\textrm{true}}$ respectively, can be evaluated by different means, for example using the L1 norm, the Hellinger distance, the Bhattacharyya distance, the total variation distance between probabilities, the Kullback-Leibler (KL) or Renyi divergences, among others. The KL divergence is used in the present work. KL is not symmetric and never negative and was used in many machine learning algorithms based on probability theory and on artificial intelligence. In information theory, KL is the amount of information lost when one distribution (for example, depending on parameters) is used to approximate another (for example, related to a ground truth). We compute the KL divergence as follows:
@@ 2.1 Case 1: A prior on onsets is available
In this expression, ${\mathbf{p}}_{\textrm{est}}$ is dependent on $S$, the subset of features used for clustering, and $K$ the number of clusters. The KL thus comprises two terms. First, a cross entropy loss
@@
and second, Shannon’s entropy
@@
2.2 Case 2: Without a prior on onsets
Even though some initial onsets could be obtained numerically (by simulations) or experimentally (for example using infrared thermography or strain sensors), it is more frequent, in practice, to be in the situation where there is no prior on onsets of clusters. Therefore, one approach to set ${\mathbf{p}}_{\textrm{true}}$ without prior is to make use of the principle of maximum entropy, since ${\mathbf{p}}_{\textrm{true}}$ is a distribution. This ensures that unconscious arbitrary assumptions are not made, and the introduction of biases extrinsic to the data is avoided [25, 15]. Therefore, in the absence of a prior on onsets, among all possible distributions of onsets for ${\mathbf{p}}_{\textrm{true}}$, we should select the one with the largest entropy that corresponds to the uniform distribution.
When using the uniform distribution of onsets for ${\mathbf{p}}_{\textrm{true}}$ in CE (Eq. 4), we obtain